# Yujie Zhou1,4∗ Pengyang Ling2,4∗ Jiazi Bu1,4∗ Yibin Wang3,5 Yuhang Zang4Jiaqi Wang4,5†Li Niu1†Guangtao Zhai1 1

### v1

@@ Abstract
The integration of online reinforcement learning (RL) into diffusion and flow models has recently emerged as a promising approach for aligning generative models with human preferences. Stochastic sampling via Stochastic Differential Equations (SDE) is employed during the denoising process to generate diverse denoising directions for RL exploration. While existing methods effectively explore potential high-value samples, they suffer from sub-optimal preference alignment due to sparse and narrow reward signals. To address these challenges, we propose a novel Granular-GRPO ($\text{G}^{2}$RPO ) framework that achieves precise and comprehensive reward assessments of sampling directions in reinforcement learning of flow models. Specifically, a Singular Stochastic Sampling strategy is introduced to support step-wise stochastic exploration while enforcing a high correlation between the reward and the injected noise, thereby facilitating a faithful reward for each SDE perturbation. Concurrently, to eliminate the bias inherent in fixed-granularity denoising, we introduce a Multi-Granularity Advantage Integration module that aggregates advantages computed at multiple diffusion scales, producing a more comprehensive and robust evaluation of the sampling directions. Experiments conducted on various reward models, including both in-domain and out-of-domain evaluations, demonstrate that our $\text{G}^{2}$RPO significantly outperforms existing flow-based GRPO baselines, highlighting its effectiveness and robustness.
11footnotetext: Equal contribution. Corresponding author.
@@ 1 Introduction
Figure 1: Comparison between our $\text{G}^{2}$RPOand existing studies. (a) $\text{G}^{2}$RPOsignificantly outperforms DanceGRPO in reward scores (HPS-v2.1 in this figure). (b) Sampling strategy comparison. $\text{G}^{2}$RPOacquire a dense reward by confining stochasticity to individual sampling steps. (c) Sampling grain comparison. $\text{G}^{2}$RPOachieves a comprehensive evaluation of each sampling direction by integrating advantages from multi-granularity ODE denoising. (d) Visual Comparison. Compared to the baseline method, the images generated by $\text{G}^{2}$RPOare more aligned with human preferences.
@@
To address these limitations, we propose Granular-GRPO ($\text{G}^{2}$RPO), a novel online reinforcement learning framework specifically designed for precise and comprehensive reward signals. First, mirroring the sparse-reward problem (Hare, 2019; Liang et al., 2024) that plagues RLHF, the reward signal in the SDE sampling process is delivered only after an entire sequence of decisions. This long delay undermines the credit-assignment chain, preventing the linking of the terminal reward with any specific earlier action and thereby inducing sluggish, unstable learning. Therefore, we propose a simple yet effective sampling strategy, termed Singular Stochastic Sampling. As illustrated in Fig 1 (b), this strategy applies the SDE formulation at a single time step to generate a group of denoising directions, while employing deterministic ODE sampling for all other steps. By concentrating stochasticity at one specific step, the proposed method establishes a strong correlation between the reward signal and the injected noise, enabling stable model optimization. Secondly, we propose a Multi-Granularity Advantage Integration (MGAI) module. As depicted in Fig. 1(c), instead of binding each denoising direction to a fixed subsequent denoising granularity, the denoising directions in the same group are assigned to a spectrum of denoising steps, producing images with different granularities. The corresponding reward signals of these images are then fused into a unified advantage estimate, yielding a comprehensive evaluation of the current state’s value.
With the support of the Singular Stochastic Sampling strategy and the Multi-Granularity Advantage Integration module, $\text{G}^{2}$RPOcan provide a more precise and comprehensive reward signal, thereby enhancing the upper limit of the GRPO model training. As shown in Fig 1 (a), our reward curves exhibit stable and significant improvements over the baseline during training. Additionally, Fig 1 (d) illustrates the images generated by $\text{G}^{2}$RPO, highlighting its advantages in text prompt adherence and detail fidelity.
Our contributions can be summarized as follows: (1) Granular-GRPO: A novel flow-based GRPO framework designed to provide a precise and comprehensive evaluation of the denoising directions sampled by the SDE, thereby improving the precision of model optimization. (2) Singular Stochastic Sampling: A sampling strategy confines stochasticity to individual sampling steps, addressing the sparse reward issue associated with long-range stochasticity injection. (3) Multi-Granularity Advantage Integration: A module integrates the advantages of multi-granularity denoised images and enables a comprehensive evaluation of each sampling direction. (4) Superior Performance: Extensive experiments across various reward models demonstrate that our $\text{G}^{2}$RPOsignificantly outperforms existing baselines, demonstrating its effectiveness and robustness.
@@ 4 Granular-GRPO
As an online RL algorithm, flow-based GRPO methods utilize SDE to sample a group of denoising directions for optimization. A core issue underlying this paradigm is to obtain precise and comprehensive assessments of each sampling direction. To this end, we introduce the $\text{G}^{2}$RPOframework to (i) confine stochasticity to individual steps (Singular Stochastic Sampling) for more precise reward signals, and (ii) integrate the advantages derived from multi-granularity denoising results (Multi-Granularity Advantage Integration) to acquire a more comprehensive evaluation, as shown in Fig. 2.
Figure 2: Overview of $\text{G}^{2}$RPO. Given a text prompt and an initial noise, our Singular Stochastic Sampling strategy employs SDE sampling solely at a single step and samples a group of distinct denoising directions. Then, the Multi-Granularity Advantage Integration module executes multi-granularity ODE denoising for each direction and integrates the advantages to produce a comprehensive evaluation for each sampling direction. For simplicity, the figure shows one coarse-grained path (denoted as $c$) and one fine-grained path (denoted as $f$).
@@ 4.2 Multi-Granularity Advantage Integration
Algorithm 1 $\text{G}^{2}$RPOTraining Process
@@ 5.1 Implementation Details
Datasets and Backbone. Following DanceGRPO (Xue et al., 2025) and MixGRPO (Li et al., 2025), we evaluate our $\text{G}^{2}$RPOusing the HPSv2 (Wu et al., 2023) dataset. It contains 103,700 text prompts for training and 400 diverse prompts for testing. The text-to-image model employed for reinforcement learning is Flux.1-dev (Labs, 2024), a leading flow model in the community.
Evaluation Metrics. To evaluate the effectiveness and robustness of our $\text{G}^{2}$RPO, multiple reward models are applied as evaluation metrics. These reward models assess the alignment between generated images and human preferences from multiple dimensions. Specifically, HPS-v2.1 (Wu et al., 2023), CLIP Score (Radford et al., 2021), and Pick Score (Kirstain et al., 2023) collectively assess semantic alignment and visual coherence. Image Reward (Xu et al., 2023) focuses on visual quality and aesthetic appeal, while Unified Reward (Wang et al., 2025) is the SOTA unified reward model that comprehensively evaluates both alignment with the caption and overall image quality.
@@ 5.2 Main Results
Quantitative Evaluation. As shown in Tab. 1, DanceGRPO and MixGRPO serve as the baselines for comparison with our $\text{G}^{2}$RPO. Additionally, the performance of employing the Singular Stochastic Sampling strategy without multi-granularity denoising ($\text{G}^{2}$RPOw/o MGAI) is also proposed. It can be observed that when HPS-v2.1 is used solely as the training reward model, our Singular Stochastic Sampling achieves a relative improvement of 6.52% compared to DanceGRPO. This indicates that constraining the stochasticity to a single step yields a precise reward signal, which in turn provides a more faithful optimization signal and enables GRPO to enhance the optimization ceiling. However, as demonstrated by DanceGRPO, optimizing solely with HPS-v2.1 can induce model hacking, which in turn compromises other out-of-domain evaluation metrics. Secondly, for the setting of multi-reward (HPS-v2.1 and CLIP Score) optimization, the results indicate that $\text{G}^{2}$RPOattains superior performance across both in-domain and out-of-domain rewards. Under the multi-granularity denoising condition, groups of images at different granularities provide a more comprehensive evaluation for the SDE sampling direction. This multi-granularity paradigm also allows for more flexible adaptation to the preferences of various reward models, thereby achieving significant improvements in various out-of-domain dimensions.
@@
$\text{G}^{2}$RPOw/o MGAI 0.372 0.395 0.234 1.421 3.688
$\text{G}^{2}$RPO 0.376 0.406 0.235 1.483 3.783
Qualitative Comparison. Fig. 4 presents the qualitative comparison among the original Flux.1-dev, DanceGRPO, MixGRPO, and our proposed $\text{G}^{2}$RPO. It can be observed that $\text{G}^{2}$RPOprovides enhanced detail fidelity and improves the consistency with the text prompt, achieving superior alignment with human preferences. For instance, in the second column, our $\text{G}^{2}$RPOfaithfully captures the specified expressions and even the nuances of the chess pieces as described in the prompt, delivering finer details and higher visual quality. Moreover, in the “poster” case depicted in the last column, $\text{G}^{2}$RPOnot only adheres to the spatial requirement of a clear left-right demarcation but also renders the reflections of the trees with remarkable clarity. Additionally, the overall style of the image generated by $\text{G}^{2}$RPOis more consistent with the aesthetic demands of poster design.
Figure 4: Qualitative Results.Comparison with existing flow-based GRPO methods, in which our $\text{G}^{2}$RPOdemonstrates superior performance in human preference alignment.
@@ 5.4 Further exploration of varying inference steps
In the main experiment Tab. 1, we maintained the same inference settings as DanceGRPO and MixGRPO, generating images with a resolution of $1024\times 1024$ in $50$ steps. Notably, our MGAI module, which evaluates denoising samples of different granularities in a mixed manner, provides a more comprehensive assessment. With the assistance of this module, $\text{G}^{2}$RPOexhibits stronger robustness to varying denoising step configurations. As shown in Tab. 3, all flow-based GRPO methods are jointly trained with HPS-v2.1 and CLIP. When the total inference timesteps are reduced to $20$ or even $10$ steps, $\text{G}^{2}$RPOstill achieved significant performance improvements across various in-domain and out-of-domain evaluation reward models.
@@ 6 Conclusion
This paper addresses the critical limitations of precisely evaluating the quality of denoising directions sampled by Flow-based GRPO for human preference alignment. We introduce $\text{G}^{2}$RPO, a novel online RL framework that precisely localizes stochasticity to a single step within the denoising process and provides a comprehensive evaluation of SDE denoising directions by integrating the advantages derived from images at different denoising granularities. This innovative design enables the provision of dense, precise reward signals, thereby fundamentally improving optimization accuracy and leading to a more robust and higher-quality alignment. Our extensive experiments consistently demonstrate that $\text{G}^{2}$RPOachieves superior performance across diverse reward conditions, marking a significant advancement in aligning generative models with human preferences.

### v2

# G2RPO: Granular GRPO for Precise Reward in Flow Models

@@ Abstract
The integration of online reinforcement learning (RL) into diffusion and flow models has recently emerged as a promising approach for aligning generative models with human preferences. Stochastic sampling via Stochastic Differential Equations (SDE) is employed during the denoising process to generate diverse denoising directions for RL exploration. While existing methods effectively explore potential high-value samples, they suffer from sub-optimal preference alignment due to sparse and narrow reward signals. To address these challenges, we propose a novel Granular-GRPO (G2RPO) framework that achieves precise and comprehensive reward assessments of sampling directions in reinforcement learning of flow models. Specifically, a Singular Stochastic Sampling strategy is introduced to support step-wise stochastic exploration while enforcing a high correlation between the reward and the injected noise, thereby facilitating a faithful reward for each SDE perturbation. Concurrently, to eliminate the bias inherent in fixed-granularity denoising, we introduce a Multi-Granularity Advantage Integration module that aggregates advantages computed at multiple diffusion scales, producing a more comprehensive and robust evaluation of the sampling directions. Experiments conducted on various reward models, including both in-domain and out-of-domain evaluations, demonstrate that our G2RPO significantly outperforms existing flow-based GRPO baselines, highlighting its effectiveness and robustness.
11footnotetext: Equal contribution Corresponding author
@@ 1 Introduction
Figure 1: Comparison between our G2RPO and existing studies. (a) G2RPO significantly outperforms DanceGRPO in reward scores (HPS-v2.1 in this figure). (b) Sampling strategy comparison. G2RPO acquire a dense reward by confining stochasticity to individual sampling steps. (c) Sampling grain comparison. G2RPO achieves a comprehensive evaluation of each sampling direction by integrating advantages from multi-granularity ODE denoising. (d) Visual Comparison. Compared to the baseline method, the images generated by G2RPO are more aligned with human preferences.
@@
To address these limitations, we propose Granular-GRPO (G2RPO), a novel online reinforcement learning framework specifically designed for precise and comprehensive reward signals. First, mirroring the sparse-reward problem (Hare, 2019; Liang et al., 2024) that plagues RLHF, the reward signal in the SDE sampling process is delivered only after an entire sequence of decisions. This long delay undermines the credit-assignment chain, preventing the linking of the terminal reward with any specific earlier action and thereby inducing sluggish, unstable learning. Therefore, we propose a simple yet effective sampling strategy, termed Singular Stochastic Sampling. As illustrated in Fig 1 (b), this strategy applies the SDE formulation at a single time step to generate a group of denoising directions, while employing deterministic ODE sampling for all other steps. By concentrating stochasticity at one specific step, the proposed method establishes a strong correlation between the reward signal and the injected noise, enabling stable model optimization. Secondly, we propose a Multi-Granularity Advantage Integration (MGAI) module. As depicted in Fig. 1(c), instead of binding each denoising direction to a fixed subsequent denoising granularity, the denoising directions in the same group are assigned to a spectrum of denoising steps, producing images with different granularities. The corresponding reward signals of these images are then fused into a unified advantage estimate, yielding a comprehensive evaluation of the current state’s value.
With the support of the Singular Stochastic Sampling strategy and the Multi-Granularity Advantage Integration module, G2RPO can provide a more precise and comprehensive reward signal, thereby enhancing the upper limit of the GRPO model training. As shown in Fig 1 (a), our reward curves exhibit stable and significant improvements over the baseline during training. Additionally, Fig 1 (d) illustrates the images generated by G2RPO, highlighting its advantages in text prompt adherence and detail fidelity.
Our contributions can be summarized as follows: (1) Granular-GRPO: A novel flow-based GRPO framework designed to provide a precise and comprehensive evaluation of the denoising directions sampled by the SDE, thereby improving the precision of model optimization. (2) Singular Stochastic Sampling: A sampling strategy confines stochasticity to individual sampling steps, addressing the sparse reward issue associated with long-range stochasticity injection. (3) Multi-Granularity Advantage Integration: A module integrates the advantages of multi-granularity denoised images and enables a comprehensive evaluation of each sampling direction. (4) Superior Performance: Extensive experiments across various reward models demonstrate that our G2RPO significantly outperforms existing baselines, demonstrating its effectiveness and robustness.
@@ 4 Granular-GRPO
As an online RL algorithm, flow-based GRPO methods utilize SDE to sample a group of denoising directions for optimization. A core issue underlying this paradigm is to obtain precise and comprehensive assessments of each sampling direction. To this end, we introduce the G2RPO framework to (i) confine stochasticity to individual steps (Singular Stochastic Sampling) for more precise reward signals, and (ii) integrate the advantages derived from multi-granularity denoising results (Multi-Granularity Advantage Integration) to acquire a more comprehensive evaluation, as shown in Fig. 2.
Figure 2: Overview of G2RPO. Given a text prompt and an initial noise, our Singular Stochastic Sampling strategy employs SDE sampling solely at a single step and samples a group of distinct denoising directions. Then, the Multi-Granularity Advantage Integration module executes multi-granularity ODE denoising for each direction and integrates the advantages to produce a comprehensive evaluation for each sampling direction. For simplicity, the figure shows one coarse-grained path (denoted as $c$) and one fine-grained path (denoted as $f$).
@@ 4.2 Multi-Granularity Advantage Integration
Algorithm 1 G2RPO Training Process
@@ 5.1 Implementation Details
Datasets and Backbone. Following DanceGRPO (Xue et al., 2025) and MixGRPO (Li et al., 2025), we evaluate our G2RPO using the HPSv2 (Wu et al., 2023) dataset. It contains 103,700 text prompts for training and 400 diverse prompts for testing. The text-to-image model employed for reinforcement learning is Flux.1-dev (Labs, 2024), a leading flow model in the community.
Evaluation Metrics. To evaluate the effectiveness and robustness of our G2RPO, multiple reward models are applied as evaluation metrics. These reward models assess the alignment between generated images and human preferences from multiple dimensions. Specifically, HPS-v2.1 (Wu et al., 2023), CLIP Score (Radford et al., 2021), and Pick Score (Kirstain et al., 2023) collectively assess semantic alignment and visual coherence. Image Reward (Xu et al., 2023) focuses on visual quality and aesthetic appeal, while Unified Reward (Wang et al., 2025) is the SOTA unified reward model that comprehensively evaluates both alignment with the caption and overall image quality.
@@ 5.2 Main Results
Quantitative Evaluation. As shown in Tab. 1, DanceGRPO and MixGRPO serve as the baselines for comparison with our G2RPO. Additionally, the performance of employing the Singular Stochastic Sampling strategy without multi-granularity denoising (G2RPO w/o MGAI) is also proposed. It can be observed that when HPS-v2.1 is used solely as the training reward model, our Singular Stochastic Sampling achieves a relative improvement of 6.52% compared to DanceGRPO. This indicates that constraining the stochasticity to a single step yields a precise reward signal, which in turn provides a more faithful optimization signal and enables GRPO to enhance the optimization ceiling. However, as demonstrated by DanceGRPO, optimizing solely with HPS-v2.1 can induce model hacking, which in turn compromises other out-of-domain evaluation metrics. Secondly, for the setting of multi-reward (HPS-v2.1 and CLIP Score) optimization, the results indicate that G2RPO attains superior performance across both in-domain and out-of-domain rewards. Under the multi-granularity denoising condition, groups of images at different granularities provide a more comprehensive evaluation for the SDE sampling direction. This multi-granularity paradigm also allows for more flexible adaptation to the preferences of various reward models, thereby achieving significant improvements in various out-of-domain dimensions.
@@
G2RPO w/o MGAI 0.372 0.395 0.234 1.421 3.688
G2RPO 0.376 0.406 0.235 1.483 3.783
Qualitative Comparison. Fig. 4 presents the qualitative comparison among the original Flux.1-dev, DanceGRPO, MixGRPO, and our proposed G2RPO. It can be observed that G2RPO provides enhanced detail fidelity and improves the consistency with the text prompt, achieving superior alignment with human preferences. For instance, in the second column, our G2RPO faithfully captures the specified expressions and even the nuances of the chess pieces as described in the prompt, delivering finer details and higher visual quality. Moreover, in the “poster” case depicted in the last column, G2RPO not only adheres to the spatial requirement of a clear left-right demarcation but also renders the reflections of the trees with remarkable clarity. Additionally, the overall style of the image generated by G2RPO is more consistent with the aesthetic demands of poster design.
Figure 4: Qualitative Results.Comparison with existing flow-based GRPO methods, in which our G2RPO demonstrates superior performance in human preference alignment.
@@ 5.4 Further exploration of varying inference steps
In the main experiment Tab. 1, we maintained the same inference settings as DanceGRPO and MixGRPO, generating images with a resolution of $1024\times 1024$ in $50$ steps. Notably, our MGAI module, which evaluates denoising samples of different granularities in a mixed manner, provides a more comprehensive assessment. With the assistance of this module, G2RPO exhibits stronger robustness to varying denoising step configurations. As shown in Tab. 3, all flow-based GRPO methods are jointly trained with HPS-v2.1 and CLIP. When the total inference timesteps are reduced to $20$ or even $10$ steps, G2RPO still achieved significant performance improvements across various in-domain and out-of-domain evaluation reward models.
@@ 6 Conclusion
This paper addresses the critical limitations of precisely evaluating the quality of denoising directions sampled by Flow-based GRPO for human preference alignment. We introduce G2RPO, a novel online RL framework that precisely localizes stochasticity to a single step within the denoising process and provides a comprehensive evaluation of SDE denoising directions by integrating the advantages derived from images at different denoising granularities. This innovative design enables the provision of dense, precise reward signals, thereby fundamentally improving optimization accuracy and leading to a more robust and higher-quality alignment. Our extensive experiments consistently demonstrate that G2RPO achieves superior performance across diverse reward conditions, marking a significant advancement in aligning generative models with human preferences.